### v1

# Task-Agnostic Machine Learning-Assisted Inference

Abstract
Machine learning (ML) is playing an increasingly important role in scientific research. In conjunction with classical statistical approaches, ML-assisted analytical strategies have shown great promise in accelerating research findings. This has also opened up a whole new field of methodological research focusing on integrative approaches that leverage both ML and statistics to tackle data science challenges. One type of study that has quickly gained popularity employs ML to predict unobserved outcomes in massive samples and then uses the predicted outcomes in downstream statistical inference. However, existing methods designed to ensure the validity of this type of post-prediction inference are limited to very basic tasks such as linear regression analysis. This is because any extension of these approaches to new, more sophisticated statistical tasks requires task-specific algebraic derivations and software implementations, which ignores the massive library of existing software tools already developed for complex inference tasks and severely constrains the scope of post-prediction inference in real applications. To address this challenge, we propose a novel statistical framework for task-agnostic ML-assisted inference. It provides a post-prediction inference solution that can be easily plugged into almost any established data analysis routine. It delivers valid and efficient inference that is robust to arbitrary choices of ML models, while allowing nearly all existing analytical frameworks to be incorporated into the analysis of ML-predicted outcomes. Through extensive experiments, we showcase the validity, versatility, and superiority of our method compared to existing approaches.
1 Introduction
Leveraging machine learning (ML) techniques to enhance and accelerate research has become increasingly popular in many scientific disciplines (Wang et al.,, 2023). For example, sophisticated deep learning frameworks have achieved remarkable success in predicting protein structure and interactions, which has the potential to significantly speed up the research process, save costs, and revolutionize the field of structural biology (Jumper et al.,, 2021; Abramson et al.,, 2024; Ahdritz et al.,, 2024). However, several recent studies have pointed out that statistical inference using ML-predicted outcomes may lead to invalid scientific discoveries due to the lack of consideration of ML prediction uncertainty in traditional statistical approaches (Wang et al.,, 2020; Angelopoulos et al., 2023a, ; Miao et al., 2023b, ). To address this, researchers have introduced methods that couple ML prediction on large samples and a small set of individuals with measured outcome data to ensure the validity of ML-assisted statistical inference (Wang et al.,, 2020; Angelopoulos et al., 2023a, ; Miao et al., 2023b, ).
Despite these advances, current ML-assisted inference methods can only address basic statistical tasks, including mean estimation, quantile estimation, and linear and logistic regression (Angelopoulos et al., 2023a, ; Miao et al., 2023b, ). While the same mathematical principle behind existing ML-assisted inference methods can be generalized to a broader class of M-estimation problems, specific algebraic derivations and computational implementations are required for each new statistical task. Moreover, many tasks, such as the Wilcoxon rank-sum test, do not fit into the M-estimation framework. These issues pose significant challenges to the broad application of ML-assisted inference across various scientific domains.
Historically, the statistics field has encountered similar challenges before the advent of resampling-based methods (Efron and Tibshirani,, 1994). It used to require task-specific derivation and implementation in order to obtain the standard error of any new estimator. This process is remarkably similar to the current state of ML-assisted inference, where every new task requires researchers to go through these non-trivial steps. However, with resampling-based inference, the need to manually derive standard errors was reduced. Instead, resampling methods can be universally applied to many estimation problems and easily obtain standard errors (Efron,, 1982, 1992; Efron and Tibshirani,, 1994). Inspired by this, we seek a universal approach that can incorporate ML-predicted outcomes into any existing analysis routine while ensuring the validity of inference results.
We introduce a simple protocol named PoSt-Prediction Sumstats-based (PSPS) inference. It employs existing analysis routines to generate summary statistics that are sufficient for ML-assisted inference, then produces valid inference results using these summary statistics. We summarize the key features of our method:
  • Assumption-lean and data-adaptive: It inherits the theoretical guarantees, including validity and efficiency, of the state-of-the-art methods for ML-assisted inference (Angelopoulos et al., 2023b, ; Miao et al., 2023b, ; Gan and Liang,, 2023). These guarantees are robust to the choice of ML procedure.
  • Task-agnostic and simple: Due to the fact that our method only needs summary statistics produced by existing analysis routines as its input, it can be easily adopted for a broad collection of statistical tasks that are currently unavailable or would require substantial effort to implement in the ML-assisted inference setting.
  • Privacy-preserving: It does not need any individual-level data as input. Sharing of privacy-preserving summary statistics is sufficient for real-world scientific collaboration.
The rest of this paper is organized as follows. section 2 formulates the ML-assisted inference problem and builds the intuition for our method. section 3 develops our general protocol for PSPS, establishes its asymptotic properties and optimality, and provides an extension to handle different data structures. section 4 introduces ML-assisted false discovery rate control for multiple testing correction. We then evaluate our methods on simulated and real datasets in section 5 and conclude in section 6. All proofs are deferred to appendix A.
2 Problem formulations
2.1 Setting
We focus on statistical inference problems for the parameter $\theta^{*}\equiv\theta^{*}(\mathbb{P})\in\mathbb{R}^{K}$ defined on the the joint distribution of $({\bf X},Y)\sim\mathbb{P}$, where $Y\in\mathcal{Y}$ is a scalar outcome and ${\bf X}\in\mathcal{X}$ be a $K$-dimensional vector representing features. We are interested in estimating $\theta^{*}$ using labeled data =${\cal L}=\{({\bf X}_{i},Y_{i}),i=1,\cdots,n\}\equiv({\bf X}_{\cal L},Y_{\cal L})$, unlabeled data ${\cal U}=\{{\bf X}_{i},i=n+1,\cdots,n+N\}\equiv{\bf X}_{\cal U}$, and a pre-trained ML model $\widehat{f}(\cdot):\mathcal{X}\rightarrow\mathcal{Y}$. Here, $f(\cdot)$ is a black-box function with unknown operating characteristics and can be mis-specified. We also require an algorithm $\mathcal{A}$ that inputs the labeled data ${\cal L}$ and returns a consistent and asymptotically normally distributed estimator $\widehat{\theta}$ for $\theta^{*}$. There are three common ways in the literature to estimate $\theta^{*}$.
  • Classical statistical methods apply algorithm $\mathcal{A}$ to only labeled data ${\cal L}=({\bf X}_{\cal L},Y_{\cal L})$, and returns the estimator and its corresponding standard error $[\widehat{\bm{\theta}}_{\cal L},\operatorname{Var}(\widehat{\bm{\theta}}_{\cal L})]$. Valid confidence intervals and hypothesis tests can then be constructed using the asymptotic distribution of the estimator. However, it ignores the unlabeled data and ML predictions.
  • Imputation-based methods that treat the ML prediction $\widehat{f}$ in the unlabeled data as the observed outcome, and apply algorithm $\mathcal{A}$ to ${\cal U}=({\bf X}_{\cal U},\widehat{f}_{\cal U})$. We denote the estimator and corresponding standard error as $[\widehat{\bm{\eta}}_{\cal U},\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U})]$. This has been shown to give invalid inference results and false scientific findings (Wang et al.,, 2020; Angelopoulos et al., 2023a, ; Miao et al., 2023b, ; Miao et al.,, 2024).
  • ML-assisted inference methods use both ${\cal L}$ and ${\cal U}$ as input. These approaches add a debiasing term in the loss function (or estimating equation) for M-estimation problems, thus removing biases from the imputation-based estimators and producing results that are statistically valid and universally more powerful compared to classical methods.
We first provide some intuition about ML-assisted inference and our proposed protocol using an example.
2.2 Building the intuition with mean estimation
We consider the mean estimation problem where $\theta^{*}=\mathbb{E}[Y_{i}]\equiv\mbox{arg min}_{\theta}\mathbb{E}[\frac{1}{2}\left(Y_{i}-\theta\right)^{2}]$. The classical method only takes the labeled data $Y_{\cal L}$ as input and yields an unbiased and consistent estimator for $\theta^{*}$:
$\widehat{\theta}_{\cal L}=\mbox{arg min}_{\theta}\frac{1}{n}\sum_{i=1}^{n}\frac{1}{2}\left(Y_{i}-\theta\right)^{2}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}.$ (1)
The imputation-based method only takes the unlabeled data $\widehat{f}_{\cal U}$ as input and returns
$\widehat{\eta}_{\cal U}=\mbox{arg min}_{\theta}\frac{1}{N}\sum_{i=n+1}^{n+N}\frac{1}{2}\left(\widehat{f}_{i}-\theta\right)^{2}=\frac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}_{i}.$ (2)
It is a biased and inconsistent estimator for $\mathbb{E}[Y_{i}]$ if the ML model $\widehat{f}$ is mis-specified. To address this, the ML-assisted estimator takes both labeled data $(Y_{\cal L},\widehat{f}_{\cal L})$ and unlabeled data $\widehat{f}_{\cal U}$ as input and adds a debiasing term to the loss function to rectify the bias caused by ML imputation:
$\widehat{\theta}_{\texttt{MLA}}$ $=\mbox{arg min}_{\theta}\frac{1}{2}\left\{\omega_{0}\frac{1}{N}\sum_{i=n+1}^{n+N}\left(\widehat{f}_{i}-\theta\right)^{2}-\underbrace{\left[\omega_{0}\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{f}_{i}-\theta\right)^{2}-\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}-\theta\right)^{2}\right]}_{\text{Debiasing term}}\right\}$ (3)
$=\omega_{0}\frac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}_{i}-\underbrace{\left[\omega_{0}\frac{1}{n}\sum_{i=1}^{n}\widehat{f}_{i}-\frac{1}{n}\sum_{i=1}^{n}y_{i}\right]}_{\text{Debiasing term}},$ (4)
where the added debiased loss ensures the consistency of the ML-assisted estimator, and the weight $\omega_{0}=\frac{\textnormal{Cov}_{l}[Y,\widehat{f}]/n}{\textnormal{Var}_{l}[\widehat{f}]/n+\textnormal{Var}_{u}[\widehat{f}]/N}$ ensures that the ML-assisted estimator is no less efficient than the classical estimator with arbitrary ML predictions. $\operatorname{Var}(\widehat{\theta}_{\texttt{MLA}})=\operatorname{Var}(\widehat{\theta}_{\cal L})-\frac{\operatorname{Cov}_{l}[Y,\widehat{f}]}{n\operatorname{Var}_{l}[\widehat{f}]+n^{2}\operatorname{Var}_{u}[\widehat{f}]/N}\leq\operatorname{Var}(\widehat{\theta}_{\cal L})$.
Our proposed method is motivated by the observation that the sufficient statistics of both the ML-assisted estimator $\widehat{\theta}_{\texttt{MLA}}$ and its variance $\operatorname{Var}(\widehat{\theta}_{\texttt{MLA}})$ are the following summary statistics:
$\widehat{\bm{\theta}}_{\texttt{ss}}=(\widehat{\theta}_{{\cal L}},\widehat{\eta}_{{\cal L}},\widehat{\eta}_{{\cal U}})\equiv(\frac{1}{n}\sum_{i=1}^{n}y_{i},\frac{1}{n}\sum_{i=1}^{n}\widehat{f}_{i},\frac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}_{i})\text{ and }\bm{\Sigma}_{\texttt{ss}}=\operatorname{Var}(\widehat{\bm{\theta}}_{\texttt{ss}})\equiv\begin{bmatrix}\textnormal{Var}_{l}[Y]/n&\textnormal{Cov}_{l}[Y,\widehat{f}]/n&0\\ \textnormal{Cov}_{l}[Y,\widehat{f}]/n&\textnormal{Var}_{l}[\widehat{f}]/n&0\\ 0&0&\textnormal{Var}_{u}[\widehat{f}]/N\\ \end{bmatrix}$ (5)
Moreover, they can be easily obtained by applying the same algorithm $\mathcal{A}$ (i.e., mean estimation in this example) to:
  • labeled data with observed outcome $\mathcal{A}(Y_{\cal L})\rightarrow[\widehat{\theta}_{{\cal L}},\operatorname{Var}(\widehat{\theta}_{{\cal L}})]=[\frac{1}{n}\sum_{i=1}^{n}y_{i},\textnormal{Var}_{l}[Y]/n]$
  • labeled data with predicted outcome $\mathcal{A}(\widehat{f}_{\cal L})\rightarrow[\widehat{\eta}_{{\cal L}},\operatorname{Var}(\widehat{\eta}_{{\cal L}})]=[\frac{1}{n}\sum_{i=1}^{n}\widehat{f}_{i},\textnormal{Var}_{l}[\widehat{f}]/n]$
  • unlabeled data with predicted outcome $\mathcal{A}(\widehat{f}_{\cal U})\rightarrow[\widehat{\eta}_{{\cal U}},\operatorname{Var}(\widehat{\eta}_{{\cal U}})]=[\frac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}_{i},\textnormal{Var}_{u}[\widehat{f}]/N]$
  • bootstrap samples of labeled data $\mathcal{A}[(Y_{\cal L},\widehat{f}_{\cal L})_{q},q=1,\dots,Q]$ for estimation of $\operatorname{Cov}(\widehat{\theta}_{{\cal L}},\widehat{\eta}_{{\cal L}})=\textnormal{Cov}_{l}[Y,\widehat{f}]/n$. Here, $(Y_{\cal L},\widehat{f}_{\cal L})_{q}$ represents the $q$-th bootstrap of labeled data.
Inputting these summary statistics for one-step debiasing $\omega_{0}\widehat{\eta}_{{\cal U}}-(\omega_{0}\widehat{\eta}_{{\cal L}}-\widehat{\theta}_{{\cal L}})$ recovers $\widehat{\theta}_{\texttt{MLA}}$.
To summarize, an algorithm for mean estimation, coupled with resampling, is sufficient for ML-assisted estimation. This inspires us to generalize this protocol for a broad range of statistical tasks. Our protocol illustrated in fig. 1 only requires three steps: 1) using the pre-trained ML model to predict the outcomes in both the labeled and unlabeled data, 2) applying the existing analysis routines to generate summary statistics, 3) employing a debiasing procedure using these summary statistics and producing statistically valid results.
Figure 1: Workflow of PSPS for Task-Agnostic ML-Assisted Inference
2.3 Related work
Our work is closely related to recent methods developed in the ML-assisted inference literature (Angelopoulos et al., 2023a, ; Miao et al., 2023b, ; Wang et al.,, 2020; Angelopoulos et al., 2023b, ; Gan and Liang,, 2023; Motwani and Witten,, 2023; Zrnic and Candès,, 2024), and is also related to methods for handling missing data (Rubin,, 1996; Tsiatis,, 2006) and semi-supervised inference (Azriel et al.,, 2022; Zhang and Bradic,, 2022; Deng et al.,, 2023; Zhang et al.,, 2019). While current ML-assisted inference methods modify the loss function or the estimating equation, our protocol works directly on the summary statistics. For simple problems such as mean estimation, current methods yield a closed-form solution to the optimization problem. However, for more general statistical tasks, there is no such closed-form solution. Current methods typically require the algebraic form of the loss function, its first and second order derivatives, and the standard error for the estimator, as well as a newly implemented optimization algorithm to obtain the estimator. We use the logistic regression problem as an example. Here, ${\bm{\theta}}^{*}=\mbox{arg min}_{{\bm{\theta}}}\mathbb{E}[-Y({\bm{\theta}}{\bf X})^{\rm T}-\psi({\bf X}{\bm{\theta}})]$ and $\psi(t)=1/(1+\exp(-t))$.
[…]
with asymptotic variance ${\bf A}^{-1}{\bf V}(\omega){\bf A}^{-1}$, where ${\bf A}=\frac{1}{N+n}(\sum_{i=1}^{n}\psi^{\prime\prime}\left({\bf X}_{i}{\bm{\theta}}\right){\bf X}_{i}^{\rm T}{\bf X}_{i}+\sum_{j=1}^{N}\psi^{\prime\prime}\left({\bf X}_{i}{\bm{\theta}}\right){\bf X}_{i}^{\rm T}{\bf X}_{i})$, ${\bf V}(\omega)=\frac{n}{N}[\omega^{2}\operatorname{Var}_{n}\left((\psi^{\prime}({\bf X}_{i}{\bm{\theta}})-\widehat{f}_{i}){\bf X}_{i}^{\rm T}\right)+\operatorname{Cov}_{N+n}\left((1-\omega)\psi^{\prime}({\bf X}_{i}{\bm{\theta}}){\bf X}_{i}^{\rm T}+(\omega\widehat{f}_{i}-Y_{i}){\bf X}_{i}^{\rm T}\right)$, and $\omega$ needs to be obtained by optimization to minimize the asymptotic variance. In contrast, our protocol simply applies logistic regression $\mathcal{A}$ to
  • labeled data with observed outcomes $({\bf X}_{\cal L},Y_{\cal L})$ to obtain $[\widehat{\bm{\theta}}_{{\cal L}},\operatorname{Var}(\widehat{\bm{\theta}}_{{\cal L}})]$
  • labeled data with predicted outcomes $({\bf X}_{\cal L},\widehat{f}_{\cal L})$ to obtain [$\widehat{\bm{\eta}}_{{\cal L}},\operatorname{Var}(\widehat{\bm{\eta}}_{{\cal L}})]$
  • unlabeled data with predicted outcomes $({\bf X}_{\cal U},\widehat{f}_{\cal U})$ to obtain $[\widehat{\bm{\eta}}_{{\cal U}},\operatorname{Var}(\widehat{\bm{\eta}}_{{\cal U}})]$
  • bootstrap of labeled data $({\bf X}_{\cal L},Y_{\cal L},\widehat{f}_{\cal L})_{q},q=1,\dots,Q$ for $\operatorname{Cov}(\widehat{\bm{\theta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal L}})$,
and returns $\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal L}}-\widehat{\bm{\theta}}_{{\cal L}})$, where $\bm{\omega}_{0}=(\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})+\frac{n}{N}\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U}))^{-1}\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$. For each new statistical task, as long as an existing analysis routine can produce an estimator that is asymptotically normally distributed, our protocol can be applied in a similar manner. Additionally, the current mathematical principles guiding ML-assisted inference are applicable solely to M-estimation (Angelopoulos et al., 2023a, ; Miao et al., 2023b, ; Angelopoulos et al., 2023b, ; Gan and Liang,, 2023; Zrnic and Candès,, 2024). However, our protocol extends beyond this limitation, addressing all estimation problems with an asymptotically normally distributed estimator.
Inference relying solely on summary statistics is widely used in the statistical genetics literature for practical reasons. Summary statistics-based methods have been developed for tasks such as variance component inference and polygenic risk prediction (Bulik-Sullivan et al., 2015b, ; Bulik-Sullivan et al., 2015a, ; Ruan et al.,, 2022; Miao et al., 2023a, ; Zhao et al.,, 2022). In contrast to our work, these applications do not leverage ML predictions, but instead focus on inference using summary statistics obtained from observed outcomes. An exception is our previous study for valid genome-wide association studies on ML-predicted outcome (Miao et al.,, 2024). But our previous work focused only on linear regression modeling with application to genome-wide association studies. The PSPS introduced in this paper aims to extend ML-assisted inference to general statistical tasks.
Our work is also related to resampling-based inference, zero augmentation, and false discovery rate (FDR) control methods. Our protocol is inspired by the core principle of resampling-based inference, which replaces algebraic derivation with computation (Efron and Tibshirani,, 1994). The main difference is that we focus on how to use ML to support inference, whereas resampling-based inference focuses on bias and variance estimation. The idea of zero augmentation has been used in augmented inverse propensity weighting estimators (Robins et al.,, 1994) and in handling unmeasured confounders (Yang and Ding,, 2019) and missing data for U-statistics (Cannings and Fan,, 2022). These estimators do not incorporate ML, which is fundamental to our work. We also adapt techniques from the FDR literature (Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001; Barber and Candès,, 2015, 2019). Our unique contribution is to use ML to support FDR control, thereby increasing its statistical power, in contrast to classical methods that rely solely on labeled data.
3 Methods
3.1 General protocol for PSPS
Building on our two examples in section 2, we formalize our protocol in fig. 1 for ML-assisted inference:
Algorithm 1 PSPS for ML-assisted inference
1:A pre-trained ML model $\widehat{f}$, labeled data ${\cal L}=({\bf X}_{\cal L},Y_{\cal L})$, unlabeled data ${\cal U}={\bf X}_{\cal U}$
2:Use the ML model $\widehat{f}$ to predict the outcome in both labeled and unlabeled data.
3:Apply the algorithm $\mathcal{A}$ in the analysis routine to
  • labeled data $({\bf X}_{\cal L},Y_{\cal L})$ and obtain $[\widehat{\bm{\theta}}_{\cal L},\operatorname{Var}(\widehat{\bm{\theta}}_{\cal L})]$
  • labeled data $({\bf X}_{\cal L},\widehat{f}_{\cal L})$ and obtain $[\widehat{\bm{\eta}}_{\cal L},\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})]$
  • unlabeled data with $({\bf X}_{\cal U},\widehat{f}_{\cal U})$ and obtain $[\widehat{\bm{\eta}}_{\cal U},\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U})]$
  • $Q$ bootstrap of labeled data $({\bf X}_{\cal L},Y_{\cal L},\widehat{f}_{\cal L})_{q},q=1,\dots,Q$ and obtain $\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$.
4:Employ one-step debiasing to the the summary statistics in step2: $\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}=\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal L}}-\widehat{{\bm{\theta}}}_{{\cal L}}),$ where $\bm{\omega}_{0}=(\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})+\frac{n}{N}\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U}))^{-1}\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ and $\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})=\operatorname{Var}(\widehat{\bm{\theta}}_{\cal L})-\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}(\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})+\frac{n}{N}\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U}))^{-1}\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$
5:ML-assisted point estimator, standard error, confidence interval, and p-value
The only requirement for our protocol are: i) algorithm $\mathcal{A}$, when applied to labeled data $(\bm{X}_{\cal L},Y_{\cal L})$, returns a consistent and asymptotically normally distributed estimator of ${\bm{\theta}}^{*}$; ii) labeled and unlabeled data are i.i.d.. Under these assumptions, the summary statistics asymptotically follows:
$n^{1/2}\left(\begin{array}[]{c}\widehat{{\bm{\theta}}}_{{\cal L}}-{\bm{\theta}}^{*}\\ \widehat{\bm{\eta}}_{{\cal L}}-\bm{\eta}\\ \widehat{\bm{\eta}}_{{\cal U}}-\bm{\eta}\\ \end{array}\right)\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left\{\left(\begin{array}[]{c}{\bf 0}_{K}\\ {\bf 0}_{K}\\ {\bf 0}_{K}\end{array}\right),\left(\begin{array}[]{ccc}{\bf V}(\widehat{\bm{\theta}}_{\cal L})&{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf 0}&{\bf 0}&\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U})\end{array}\right)\right\},$ (16)
where $\bm{\eta}\equiv\bm{\eta}(\mathbb{P}_{\widehat{f}})\in\mathbb{R}^{K}$ is defined on $({\bf X},\widehat{f})\sim\mathbb{P}_{\widehat{f}}$, ${\bf V}(\cdot)$ denotes the asymptotic variance and covariance of a estimator, and $\rho=\frac{n}{N}$. The asymptotic approximation gives ${\bf V}(\widehat{\bm{\theta}}_{\cal L})=n\operatorname{Var}(\widehat{\bm{\theta}}_{\cal L}),{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})=n\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L}),{\bf V}(\widehat{\bm{\eta}}_{\cal L})=n\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})$ and ${\bf V}(\widehat{\bm{\eta}}_{\cal U})=N\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U})$. Here, we do not require $\widehat{\bm{\eta}}_{\cal L}$ and $\widehat{\bm{\eta}}_{\cal U}$ to be consistent for $\bm{\theta}^{*}$, thus allows arbitrary ML model.
With the asymptotically multivariate normal distribution for the summary statistics, the one-step debiased estimator $\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}=\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal L}}-\widehat{{\bm{\theta}}}_{{\cal L}})$ is consistent for $\theta^{*}$ and asymptotically normally distributed. The $\alpha$-level confidence interval for coordinate $k$ can be constructed by $\mathcal{C}_{\alpha}^{\textnormal{{PSPS}}}=(\widehat{{\bm{\theta}}}_{\textnormal{{PSPS}}_{k}}\pm z_{1-\alpha/2}\sqrt{\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}})$. Hypothesis testing for $H_{0}:\theta^{*}_{k}=0$ can be performed using the test statistics $\frac{\widehat{{\bm{\theta}}}_{\textnormal{{PSPS}}_{k}}}{\sqrt{\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}})}\sim\mathcal{N}(0,1)$. It gives (two-sided) p-value $=2(1-\Phi(|\frac{\widehat{{\bm{\theta}}}_{\textnormal{{PSPS}}_{k}}}{\sqrt{\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}})}|))$, where $\Phi$ is the CDF of the standard normal distribution.
3.2 Theoretical guarantees
In this section, we examine the theoretical properties of PSPS. In what follows, $\stackrel{{\scriptstyle P}}{{\rightarrow}}$ denotes convergence in probability and $\stackrel{{\scriptstyle D}}{{\rightarrow}}$ denotes convergence in distribution.
The first result shows that our proposed estimator is consistent, asymptotically normally distributed, and uniformly better in terms of element-wise asymptotic variance compared with the classical estimator based on labeled data only.
Theorem 1.
Assuming equation (16) holds, then $\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\theta}}^{*}$, and
$n^{1/2}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}-{\bm{\theta}}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left(0,{\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})\right),$ (17)
where ${\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})={\bf V}(\widehat{\bm{\theta}}_{\cal L})-{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$.
Assume the $k$-th column of ${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ is not a zero vector and at least one of ${\bf V}(\widehat{\bm{\eta}}_{\cal L})$ and ${\bf V}(\widehat{\bm{\eta}}_{\cal U})$ are positive definite, then ${\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}\leq{\bf V}(\widehat{\bm{\theta}}_{\cal L})_{kk}$.
This result indicates a greater reduction in variance for the ML-assisted estimator is associated with larger values of ${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ and smaller values of ${\bf V}(\widehat{\bm{\eta}}_{\cal L})$, ${\bf V}(\widehat{\bm{\eta}}_{\cal U})$, and $\rho$. The variance reduction term $[{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})]_{kk}$ can also serve as a metric for selecting the optimal ML model in ML-assisted inference.
Our next result shows that three existing methods, i.e., PPI, PPI++, and POP-Inf, are asymptotically equivalent to PSPS with different weighting vector. A broader class for consistent estimator of $\bm{\theta}^{*}$ is $\widehat{{\bm{\theta}}}(\bm{\omega})=\bm{\omega}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\bm{\omega}^{\rm T}\widehat{{\bm{\theta}}}_{{\cal L}}-\widehat{\bm{\eta}}_{{\cal L}})$, where $\bm{\omega}$ is a $K\times K$ matrix. The consistency of $\widehat{{\bm{\theta}}}(\bm{\omega})$ for $\bm{\theta}^{*}$ only requires $\bm{\omega}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}\mathbf{0}$. Since $(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}\mathbf{0}$, assigning arbitrary fixed weights for will satisfy the condition. However, the choice of weights influences the efficiency of the estimator as illustrated in proposition 2 later.
Proposition 1.
Assuming equation (16) and regularity condition for the asymptotically normality of current ML-assisted estimator holds. For M-estimation problem, we have
$\widehat{{\bm{\theta}}}\left({\rm diag}(\bm{\omega}_{\textnormal{ele}}){\bf C}\right)-\widehat{{\bm{\theta}}}_{\textnormal{{POP-Inf}}}\stackrel{{\scriptstyle D}}{{\rightarrow}}{\bf 0},\quad\widehat{{\bm{\theta}}}\left({\rm diag}(\bm{\omega}_{\textnormal{tr}}){\bf C}\right)-\widehat{{\bm{\theta}}}_{\textnormal{{PPI++}}}\stackrel{{\scriptstyle D}}{{\rightarrow}}{\bf 0},\quad\widehat{{\bm{\theta}}}\left({\rm diag}({\bf 1}){\bf C}\right)-\widehat{{\bm{\theta}}}_{\textnormal{{PPI}}}\stackrel{{\scriptstyle D}}{{\rightarrow}}{\bf 0}.$
Here, $\bm{\omega}_{\textnormal{ele}}=[\omega_{\textnormal{ele},1},\dots,\omega_{\textnormal{ele},K}]^{\rm T}\in\mathbb{R}^{K}$ and $\omega_{\textnormal{ele},k}$ minimizing the $k$-th diagonal element of ${\bf V}(\widehat{{\bm{\theta}}}(\bm{\omega}))$, $\omega_{tr}$ is a scalar used to minimize the trace of ${\bf V}(\widehat{{\bm{\theta}}}(\bm{\omega}))$, and ${\bf C}$ is a matrix associated with the second derivatives of the loss function in M-estimation, with further details deferred to appendix A.
This demonstrates that for M-estimation problems, our method is asymptotically equivalent to POP-Inf, PPI++, and PPI each with the respective weights ${\rm diag}(\bm{\omega}_{\textnormal{ele}}){\bf C}$, ${\rm diag}(\bm{\omega}_{\textnormal{tr}}){\bf C}$, and ${\rm diag}({\bf 1}){\bf C}$. Therefore, PSPS can be viewed as a generalization of these existing methods.
Our third result shows that the weights used in the proposition 1 are not optimal. Instead, our choice of $\bm{\omega}_{0}$ represents the optimal smooth combination of $(\widehat{{\bm{\theta}}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal U}})$ in terms of minimizing the asymptotic variance, while still preserving consistency.
Proposition 2.
Suppose $n^{1/2}(g(\widehat{{\bm{\theta}}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal U}})-{\bm{\theta}}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}(0,\bm{\Sigma}_{g})$ and $g$ is a smooth function, then $\bm{\Sigma}_{g_{kk}}\geq\bm{\Sigma}_{\textnormal{{PSPS}}_{kk}}$
Together with proposition 1, our results demonstrate that our protocol provides a more efficient estimator compared to existing methods for the M-estimation problem. Furthermore, the applicability of our protocol is not limited to M-estimation and only requires summary statistics as input. It also indicates that in a distributed inference setting (Jordan et al.,, 2018) where individual-level data are not available, PSPS proves to be the optimal approach for combining shared summary statistics.
3.3 Extensions
We also provide several extensions to ensure the broad applicability of our proposed method.
3.3.1 Labeled data and unlabeled data are not independent
Here, we relax the assumption that the labeled data and unlabeled data are independent. When they are not independent, this can leads to the non-zero covariance between the $\widehat{\bm{\eta}}_{\cal L}$ and $\widehat{\bm{\eta}}_{\cal U}$. Consider a more broader class of summary statistics asymptotically satisfying
$n^{1/2}\left(\begin{array}[]{c}\widehat{{\bm{\theta}}}_{\cal L}-{\bm{\theta}}\\ \widehat{\bm{\eta}}_{\cal L}-\bm{\eta}\\ \widehat{\bm{\eta}}_{\cal U}-\bm{\eta}\\ \end{array}\right)\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left\{\left(\begin{array}[]{c}{\bf 0}_{K}\\ {\bf 0}_{K}\\ {\bf 0}_{K}\end{array}\right),\left(\left(\begin{array}[]{ccc}{\bf V}(\widehat{\bm{\theta}}_{\cal L})&{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})\\ {\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L})&\sqrt{\rho}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})\\ {\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})&\sqrt{\rho}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})&\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U})\end{array}\right)\right)\right\}$
We can similarly employ the one-step debiasing
$\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}^{\texttt{ no-indep}}=\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{\cal U}-\bm{\omega}_{0}(\widehat{\bm{\eta}}_{\cal L}-\widehat{{\bm{\theta}}})$ (18)
where $\bm{\omega}_{0}=({\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})-{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{\rm T}({\bf V}({\widehat{\bm{\eta}}_{\cal L}})+{\bf V}({\widehat{\bm{\eta}}_{\cal U}})-2{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{-1}$ and $\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}^{\texttt{ no-indep}})=\operatorname{Var}(\widehat{\bm{\theta}}_{\cal L})-({\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})-{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{\rm T}({\bf V}({\widehat{\bm{\eta}}_{\cal L}})+{\bf V}({\widehat{\bm{\eta}}_{\cal U}})-2{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{-1}({\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})-{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))$. The theoretical guarantees of the proposed estimator can be similarly derived by theorem 1.
3.3.2 Sensitivity analysis for distributional shift between labeled and unlabeled data
The other assumption of our approach is that the labeled and unlabeled data are identically distributed so that we can ensure $\widehat{\bm{\eta}}_{\cal L}-\widehat{\bm{\eta}}_{\cal U}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bf 0}$ and validity of PSPS results. To address the potential violation of this assumption, we introduce a sensitivity analysis with hypothesis testing for the null $H_{0}:\bm{\eta}_{{\cal L},k}=\bm{\eta}_{{\cal U},k}$ with test statistics
$\frac{\widehat{\bm{\eta}}_{{\cal L},k}-\widehat{\bm{\eta}}_{{\cal U},k}}{\sqrt{\operatorname{Var}(\widehat{\bm{\eta}}_{{\cal L},k})+\operatorname{Var}(\widehat{\bm{\eta}}_{{\cal U},k})}}\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}(0,1)$ (19)
to assess if $\bm{\eta}_{{\cal L},k}$ and $\bm{\eta}_{{\cal U},k}$ are significantly different. Here, the subscript $k$ indicates the $k$-th coordinate. We recommend using p-value $<0.1$ as evidence for heterogeneity and caution the interpretation for results of ML-assisted inference.
4 ML-assisted False Discovery Rate Control
Controlling FDR is crucial for statistical application in various scientific disciplines (Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001; Barber and Candès,, 2015). In this section, we show that our protocol can be utilized for ML-assisted FDR control, achieving greater statistical power compared to classical methods that solely rely on labeled data. We consider the linear regression model with
$\bm{Y}=\sum_{k=1}^{K}\bm{X}_{k}\beta_{k}+\bm{\epsilon},\bm{\epsilon}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\bm{I}_{n}\right)$ (20)
Here, $\bm{Y}$ is a $n$-dimensional vector, $\bm{X}_{k}$ is also a $n$-dimensional vector representing the $k$-th feature. A central question here is to select the variables associated with the outcome $\bm{Y}$ after accounting for all other variables. Formally, denote $\widehat{S}\subset\{1,2,\ldots,K\}$ as the set of selected variables. The FDR is defined as
$\text{FDR}=\mathbb{E}[\frac{\#\left\{k:\beta_{k}=0,k\in\widehat{S}\right\}}{\#\{k:k\in\widehat{S}\}\vee 1}].$ (21)
We consider two popular FDR control methods: Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg,, 1995) and knockoff (Barber and Candès,, 2015). BH procedure first computes p-values $p_{k}$ for each feature ${\bf X}_{k}$ using multiple linear regression and then applies a threshold to these p-values. Features with p-values smaller than the threshold are considered as discoveries. The knockoff method does not rely on p-values. It includes three key components: a ranking algorithm that assigns an importance metric to each variable, an augmented design that creates fake variables, and a symmetric statistic that utilizes the importance metric for each original variable to control the final FDR (Ke et al.,, 2024).
We introduce PSPS-BH and PSPS-knockoff for ML-assisted FDR control. Briefly, PSPS-BH processes the p-value from ML-assisted linear regression through the BH procedure, while PSPS-knockoff utilizes the ML-assisted debiased Lasso coefficient (Javanmard and Montanari,, 2014; Zhang and Zhang,, 2014; Liu and Rigollet,, 2019), in a ranking algorithm. These methods aim to improve the power of classical methods that employ linear regression or debiased Lasso (Zhang and Zhang,, 2014) using only labeled data. Detailed descriptions of the algorithms are provided below:
Algorithm 2 PSPS-BH for linear regression
1:Labeled data ${\cal L}=({\bf X}_{\cal L},Y_{\cal L},\widehat{f}_{\cal L})$, unlabeled data ${\cal U}=({\bf X}_{\cal U},\widehat{f}_{\cal U})$, FDR level $q\in(0,1)$.
2:Obtain the p-value $p_{k}$ for features $k=1,\cdots,K$ by ML-assisted linear regression PSPS-LR(${\cal L},{\cal U}$)
3:Sort the p-values in ascending order $p_{(1)}\leq p_{(2)}\leq\ldots\leq p_{(K)}$
4:Finds the p-value cutoff $\tau^{\mathrm{BH}}_{q}:=p_{(k)},\text{ where }k=\max\left\{k=1,\ldots,K:p_{(k)}\leq\frac{k}{K}q\right\}$
5:Discoveries $\widehat{S}=\left\{k:p_{k}\leq\tau^{\mathrm{BH}}_{q}\right\}$
Algorithm 3 PSPS-knockoff with debiased Lasso
1:Labeled data ${\cal L}=({\bf X}_{\cal L},Y_{\cal L},\widehat{f}_{\cal L})$, unlabeled data ${\cal U}=({\bf X}_{\cal U},\widehat{f}_{\cal U})$, FDR level $q\in(0,1)$.
2:Obtain the augmented labeled and unlabeled data as ~$\tilde{{\cal L}}=({\bf X}_{\cal L},\tilde{{\bf X}}_{\cal L},Y_{\cal L},\widehat{f}_{\cal L})$ and $\tilde{{\cal U}}=({\bf X}_{\cal U},\tilde{{\bf X}}_{\cal U},\widehat{f}_{\cal U})$ where $\tilde{{\bf X}}_{\cal L}\leftarrow$ knockoff-sample$(\bm{X}_{\cal L})$ and $\tilde{{\bf X}}_{\cal U}\leftarrow$ knockoff-sample$(\bm{X}_{\cal U})$.
3:Calculate the PSPS debiased Lasso coefficient $\widehat{\bm{\beta}}^{\texttt{PSPS-DLasso}}\leftarrow\texttt{PSPS-DLasso}(\tilde{{\cal L}},\tilde{{\cal U}})$
4:$W_{k}\text{ for }k=1,\cdots,K\leftarrow\texttt{knockoff-statistic}(\widehat{\bm{\beta}}^{\texttt{PSPS-DLasso}})$
5:Set the cutoff $\tau_{q}^{\textnormal{knockoff}}=\left\{t>0:\frac{\#\left\{k:W_{k}\leq-t\right\}}{\#\left\{k:W_{k}\geq t\right\}\vee 1}\leq q\right\}$
6:Discoveries $\widehat{S}=\left\{k:M_{k}>\tau_{q}^{\textnormal{knockoff}}\right\}$
Here, we employ second-order multivariate Gaussian knockoff variables for knockoff-sample and use the difference between the absolute values of the $k$-th feature and its knockoff coefficient as the knockoff-statistic. Alternative choices for these two steps can also be readily integrated into our algorithm (Ke et al.,, 2024).
We establish the asymptotic FDR control for our proposed method, with the same guarantees as the classical method that only using the labeled data.
Proposition 3.
Assuming positive regression dependence (PRDS) on the p-values (Benjamini and Yekutieli,, 2001), PSPS-BH with ML-assisted linear regression PSPS-LR asymptotically controls the FDR at or below level $qK_{0}/K$, where $K_{0}$ and $K$ is the number of true null and total hypotheses, respectively.
Proposition 4.
PSPS-knockoff with ML-assisted debiased Lasso PSPS-DLasso as the ranking algorithm asymptotically controls the FDR at level $q$.
5 Numerical experiments and real data application
5.1 Simulations
We conduct simulations to assess the finite sample performance of our method. Our objectives are to demonstrate that 1) PSPS achieves narrower confidence intervals when applied to statistical tasks already implemented in existing ML-assisted methods; 2) when applied to statistical tasks that have not been implemented for ML-assisted inference, PSPS provides confidence intervals with narrower width but correct coverage (indicating higher statistical power) compared to classical approaches rely solely on labeled data; 3) it provides accurate FDR control and achieves higher power compared to classical methods for FDR control which only use labeled data.
Tasks that have been implemented for ML-assisted inference We compare PSPS with the classical method using only labeled data, the imputation-based method using only unlabeled data, and three ML-assisted inference methods PPI, PPI++, and POP-Inf (Angelopoulos et al., 2023a, ; Angelopoulos et al., 2023b, ; Miao et al., 2023b, ) on mean estimation, linear and logistic. We defer the detailed data generating process to appendix B. In short, we generated outcome $Y_{i}$ from feature $X_{1i}$ and $X_{2i}$, and obtained the pre-trained random forest that predict $Y_{i}$ using $X_{1i}$ and $X_{2i}$. We have 500 labeled samples $(X_{1i},Y_{i},\widehat{f}_{i})$ , and unlabeled samples $(X_{1i},\widehat{f}_{i})$ ranged from 1,000 to 10,000. Our goal is to estimate the mean of $Y_{i}$, the linear and logistic regression coefficient between $Y_{i}$ and $X_{1i}$.
fig. 2a-c shows confidence interval coverage and fig. 2d-f shows confidence interval width. We find that imputation-based method fail to obtain correct coverage, while all other including PSPS have the correct the coverage. PSPS has narrower confidence interval width compared to the classical method and other approaches for ML-assisted inference.
Figure 2: Simulation for implemented tasks including mean estimation, linear and logistic regression from left to right. Panel a-c shows coverage and d-f shows confidence interval width.
Tasks that have not been implemented for ML-assisted inference Next, we consider several commonly used statistical tasks that currently lack implementations for ML-assisted inference including quantile regression (Koenker,, 2005), instrumental variable (IV) regression (Angrist and Imbens,, 1995), negative binomial (NB) regression (Hilbe,, 2011), debiased Lasso (Zhang and Zhang,, 2014), and the Wilcoxon rank-sum test (Mann and Whitney,, 1947). Similar to our previous simulations, we utilize labeled data, unlabeled data, and a pre-trained ML model. Detailed simulation settings are deferred to the appendix B. Our goal is to estimate the regression coefficient between $Y_{i}$ and $X_{1i}$ for quantile (at quantile level 0.75), IV, and NB regression, between $Y_{i}$ and high dimensional features ${\bf X}_{i}\in\mathbb{R}^{150}$ for debiased Lasso, and to perform hypothesis testing on the medians of two independent samples $Y_{i}|X_{1i}=1$ and $Y_{i}|X_{1i}=0$ using the Wilcoxon rank-sum test.
fig. 3a-d shows confidence interval coverage and fig. 3f-i shows confidence interval width for parameter estimation. fig. 3e and fig. 3j shows the type-I error and statistical power for Wilcoxon rank-sum test. We found that imputation-based method fail to obtain correct coverage or have inflated type-I error, while PSPS and classical method have the correct the coverage and type-I error control. PSPS have narrower confidence intervals width in all settings, and higher statistical power for Wilcoxon rank-sum test compared to classical methods. Confidence intervals become narrower as unlabeled sample size increases, indicating higher efficiency gain.
Figure 3: Simulation for not implemented tasks including quantile, instrumental variable (IV), negative binomial (NB) regression, debiased Lasso, and Wilcoxon rank-sum test from left to right. Panel a-e shows coverage (1 - type-error for Wilcoxon rank-sum test) and f-j shows confidence interval width (1 - power for Wilcoxon rank-sum test).
FDR control We evaluate the finite sample performance of PSPS-BH and PSPS-knockoff in controlling the FDR compared with classical and imputation-based methods as baselines. We consider low–dimensional($K<n$) and high-dimensional($K>n$) linear regressions for PSPS-BH and PSPS-knockoff, respectively. We simulate the data such that only a proportion of the features are truly associated with the outcome. The data generating process is deferred to appendix B. Our goal is to select the associated features while maintaining the target FDR level.
fig. 4a-b shows the estimated FDR and fig. 4c-d shows the statistical power for different methods. Imputation-based control method failed to control FDR in either low-dimensional or high-dimensional settings. Classical approach, PSPS-BH, and PSPS-knockoff effectively controlled in both low-dimensional and high-dimensional settings. PSPS-BH, and PSPS-knockoff achieve higher statistical power compared to the classical method.
Figure 4: Simulation for FDR control. Panel a-b shows the estimated FDR level given expected FDR. Panel c-d shows the power.
Together, these simulations demonstrate that PSPS outperforms existing methods and can be easily adopted for a variety of important tasks that have not yet been implemented in existing ML-assisted inference methods.
5.2 Identify vQTLs for bone mineral density
We employed our method to carry out ML-assisted quantile regression to identify genetic variants associated with the outcome variability (vQTL) of bone mineral density derived from dual-energy X-ray absorptiometry imaging (DXA-BMD) (Miao et al.,, 2022). DXA-BMD serves as the primary diagnostic marker for osteoporosis and fracture risk (Zheng et al.,, 2015; Carey and Delaney,, 2017). Identifying vQTL for DXA-BMD can provide insights into the biological mechanisms underlying outcome plasticity and reveal candidate genetic variants potential gene-gene and gene-environment interactions (Wang et al.,, 2019; Marderstein et al.,, 2021; Westerman et al.,, 2022; Young et al.,, 2018; Miao and Lu,, 2022). We focused on total body DXA-BMD, which integrates measurements from multiple skeletal sites. We used data from the UK Biobank (Bycroft et al.,, 2018), which includes 36,971 labeled and 319,548 unlabeled samples with 9,450,880 genetic variants after quality control. Data processing details have been reported elsewhere (Miao et al.,, 2024). We predicted DXA-BMD values in both labeled and unlabeled samples using SoftImpute (Mazumder et al.,, 2010) with 466 other variables measured in UK Biobank. Prediction in the labeled sample was implemented through cross-validation to avoid overfitting. We used the BH procedure to correct for multiple testing and considered FDR $<0.05$ as the significance threshold.
No genetic variants reached statistical significance under the classical method with only labeled data. PSPS identified 108 significant variants with FDR $<0.05$ spanning 5 independent loci, showcasing the superior statistical power of PSPS (fig. 5 and table 1). Notably, these the significant vQTL cannot be identified by linear regression (Miao et al.,, 2024), indicating the different genetic mechanisms controlling outcome levels and variability for DXA-BMD.
Figure 5: Manhattan plot of vQTLs for bone mineral density. X-axis represents chromosomes (CHR), plotted by base pair positions (BP). Each point on the plot indicates a single nucleotide polymorphism (SNP). Y-axis depicts -log10(p-values).
CHR BP SNP A1 A2 EAF BETA SE p-value FDR
1 205359339 rs12139623 T C 0.105 0.063 0.012 7.2e-08 0.033
4 14359045 rs552582509 A G 0.213 0.046 0.009 4.5e-07 0.045
7 132568586 rs79089873 A G 0.073 -0.068 0.014 5.6e-07 0.049
9 11587905 rs146938822 A G 0.022 -0.134 0.024 4.0e-08 0.033
15 47672201 rs281258 C T 0.393 -0.04 0.007 5.8e-08 0.033
Table 1: Significant vQTLs for bone mineral density. Abbreviations: CHR, Chromosome; BP, Base Pair; SNP, Single Nucleotide Polymorphism; A1, Allele 1 (Effect Allele); A2, Allele 2 (Non-effect Allele); EAF, Effect Allele Frequency; BETA, Effect Size (Beta Coefficient); SE, Standard Error; FDR, False Discovery Rate.
6 Conclusion
In this work, we introduced a simple, task-agnostic protocol for ML-assisted inference, with applications across a broad range of statistical tasks. We established consistency and asymptotic normality of the proposed estimator. Additionally, we introduce several extensions to further expand the scope of our approach, addressing challenges such as dependence and distributional shifts between labeled and labeled data, and FDR control. Through extensive simulations and applications to real data, we demonstrate the superior performance and broad applicability of our method across diverse tasks. As a future direction, it would be interesting to develop and implement a fast algorithm for resampling in our method.
The codes to implement PSPS, along with the benchmark methods, and to replicate the simulation and real data analysis, are available at https://github.com/qlu-lab/PSPS.
Acknowledgments
The authors gratefully acknowledge research support from National Institutes of Health (NIH) grant U01 HG012039, the University of Wisconsin-Madison Office of the Chancellor, and the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation (WARF).
Appendix A Proofs
A.1 Proof the Theorem 1
Proof.
Since
$n^{1/2}\left(\begin{array}[]{c}\widehat{\bm{\theta}}_{{\cal L}}-\bm{\theta}\\ \widehat{\bm{\eta}}_{{\cal L}}-\bm{\eta}\\ \widehat{\bm{\eta}}_{{\cal U}}-\bm{\eta}\\ \end{array}\right)\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left\{\left(\begin{array}[]{c}{\bf 0}_{K}\\ {\bf 0}_{K}\\ {\bf 0}_{K}\end{array}\right),\left(\begin{array}[]{ccc}{\bf V}(\widehat{\bm{\theta}}_{\cal L})&{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf 0}&{\bf 0}&\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U})\end{array}\right)\right\},$ (31)
$\widehat{\bm{\theta}}_{{\cal L}}\stackrel{{\scriptstyle P}}{{\rightarrow}}\bm{\theta}$ and $\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bf 0}$. Given fixed weights $\bm{\omega}_{0}=({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$, $\bm{\omega}_{0}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bf 0}$.
By Slutsky’s theorem,
$\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}=\widehat{\bm{\theta}}_{{\cal L}}+\bm{\omega}_{0}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}\bm{\theta},$ (32)
which completes the proof of consistency.
By multivariate delta methods, denoting $h([{\bf x},\mathbf{y},{\bf z}]^{\rm T})={\bf x}+\bm{\omega}_{0}^{\rm T}({\bf z}-\mathbf{y})$, we have $\nabla h([{\bf x},\mathbf{y},{\bf z}]^{\rm T})=[{\bf 1},-\bm{\omega}_{0},\bm{\omega}_{0}]$, therefore
$n^{1/2}h[\left(\begin{array}[]{c}\widehat{\bm{\theta}}_{{\cal L}}-\bm{\theta}^{*}\\ \widehat{\bm{\eta}}_{{\cal L}}-\bm{\eta}\\ \widehat{\bm{\eta}}_{{\cal U}}-\bm{\eta}\\ \end{array}\right)]$ $=n^{1/2}[\widehat{\bm{\theta}}_{{\cal L}}+\bm{\omega}_{0}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})]$ (36)
$\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}(\bm{\theta},{\bf V}(\widehat{\bm{\theta}}_{\cal L})-{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})),$ (37)
which completes the proof of asymptotic normality.
Denote ${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})_{:,k}$ as the $k$-th column of ${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$, the $k$-th diagonal element of ${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ is a quadratic form
${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})_{:,k}^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})_{:,k}.$ (38)
Here, by our assumption, $({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}$ is a positive definite matrix. Therefore, quadratic form ${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})_{:,k}^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})_{:,k}$ is non-negative and is zero if only all elements of ${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})_{:,k}$ is zero, which completes the proof of element-wise variance reduction. ∎
A.2 Proof the Proposition 1
Proof.
Given
$n^{1/2}\left(\begin{array}[]{c}\widehat{\bm{\theta}}_{{\cal L}}-\bm{\theta}^{*}\\ \widehat{\bm{\eta}}_{{\cal L}}-\bm{\eta}\\ \widehat{\bm{\eta}}_{{\cal U}}-\bm{\eta}\\ \end{array}\right)\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left\{\left(\begin{array}[]{c}{\bf 0}_{K}\\ {\bf 0}_{K}\\ {\bf 0}_{K}\end{array}\right),\left(\begin{array}[]{ccc}{\bf V}(\widehat{\bm{\theta}}_{\cal L})&{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf 0}&{\bf 0}&\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U})\end{array}\right)\right\},$ (48)
the asymptotic variance of $\widehat{\bm{\theta}}(\bm{\omega})=\widehat{\bm{\theta}}(\bm{\omega})=\widehat{\bm{\theta}}_{{\cal L}}+\bm{\omega}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})$ is
${\bf V}(\widehat{\bm{\theta}}(\bm{\omega}))={\bf V}(\widehat{\bm{\theta}}_{\cal L})+\bm{\omega}^{\rm T}{\bf V}(\widehat{\bm{\eta}}_{\cal L})\bm{\omega}+\bm{\omega}^{\rm T}\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U})\bm{\omega}-2\bm{\omega}^{\rm T}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ (49)
We first define the M-estimation (Z-estimation) problem. The goal is to estimate a $K$-dimensional parameter $\bm{\theta}^{*}$ defined through an estimating equation
$\mathbb{E}\{\bm{\psi}(Y,\mathbf{X};\bm{\theta})\}=\mathbf{0},$ (50)
where $\bm{\psi}(\cdot,\cdot;\bm{\theta})$ is a user-defined function. By the theory of Z-estimation and recent paper on ML-assisted inference (Miao et al., 2023b, ), we have ${\bf V}(\widehat{\bm{\theta}}_{\cal L})={\bf A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{1}{\bf A}_{\bm{\theta}^{*}}^{-1},{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})={\bf A}_{\bm{\eta}}^{-1}{\bf M}_{4}{\bf A}_{\bm{\theta}^{*}}^{-1},{\bf V}(\widehat{\bm{\eta}}_{\cal L})={\bf A}_{\bm{\eta}}^{-1}{\bf M}_{2}{\bf A}_{\bm{\eta}}^{-1}$, and ${\bf V}(\widehat{\bm{\eta}}_{\cal U})={\bf A}_{\bm{\eta}}^{-1}{\bf M}_{3}{\bf A}_{\bm{\eta}}^{-1}$. Here, ${\bf M}_{1}=\textnormal{Var}_{l}[\bm{\psi}(Y,{\bf X};\bm{\theta}^{*})],{\bf M}_{2}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f},{\bf X};\bm{\theta}^{*})],{\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f},{\bf X};\bm{\theta}^{*})],{\bf M}_{4}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};\bm{\theta}^{*}),\bm{\psi}(\widehat{f},{\bf X};\bm{\theta}^{*})],{\bf A}_{\bm{\theta}^{*}}=\mathbb{E}[\partial\bm{\psi}(Y,{\bf X};\bm{\theta}^{*})/\partial{\bm{\theta}}],{\bf A}_{\bm{\eta}}=\mathbb{E}[\partial\bm{\psi}(\widehat{f},{\bf X};\bm{\eta})/\partial\bm{\eta}]$, and $\rho=\dfrac{n}{N}$.
Rewritten ${\bf V}(\widehat{\bm{\theta}}(\bm{\omega}))$ using the above notation, we have
${\bf V}(\widehat{\bm{\theta}}(\bm{\omega}))$ $={\bf A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{1}{\bf A}_{\bm{\theta}^{*}}^{-1}+\bm{\omega}^{\rm T}{\bf A}_{\bm{\eta}}^{-1}{\bf M}_{2}{\bf A}_{\bm{\eta}}^{-1}\bm{\omega}+\rho\bm{\omega}^{\rm T}{\bf A}_{\bm{\eta}}^{-1}{\bf M}_{3}{\bf A}_{\bm{\eta}}^{-1}\bm{\omega}-2\bm{\omega}^{\rm T}{\bf A}_{\bm{\eta}}^{-1}{\bf M}_{4}{\bf A}_{\bm{\theta}^{*}}^{-1}$ (51)
$={\bf A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{1}{\bf A}_{\bm{\theta}^{*}}^{-1}+\bm{\omega}^{\rm T}{\bf A}_{\bm{\eta}}^{-1}({\bf M}_{2}+\rho{\bf M}_{3}){\bf A}_{\bm{\eta}}^{-1}\bm{\omega}-2\bm{\omega}^{\rm T}{\bf A}_{\bm{\eta}}^{-1}{\bf M}_{4}{\bf A}_{\bm{\theta}^{*}}^{-1}$ (52)
Plug in
$\bm{\omega}=\bm{\omega}_{0}$ $=(\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})+\frac{n}{N}\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U}))^{-1}\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ (53)
$=({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ (54)
$=({\bf A}_{\bm{\eta}}^{-1}{\bf M}_{2}{\bf A}_{\bm{\eta}}^{-1}+\rho{\bf A}_{\bm{\eta}}^{-1}{\bf M}_{3}{\bf A}_{\bm{\eta}}^{-1})^{-1}{\bf A}_{\bm{\eta}}^{-1}{\bf M}_{4}{\bf A}_{\bm{\theta}^{*}}^{-1}$ (55)
$={\bf A}_{\bm{\eta}}({\bf M}_{2}+\rho{\bf M}_{3})^{-1}{\bf M}_{4}{\bf A}_{\bm{\theta}^{*}}^{-1}$ (56)
into the equation above, we have
${\bf V}(\widehat{\bm{\theta}}(\bm{\omega}_{0}))$ $={\bf A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{1}{\bf A}_{\bm{\theta}^{*}}^{-1}-{\bf A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{4}^{\rm T}{\bf A}_{\bm{\eta}}^{-1}[{\bf A}_{\bm{\eta}}({\bf M}_{2}+\rho{\bf M}_{3})^{-1}{\bf A}_{\bm{\eta}}]{\bf A}_{\bm{\eta}}^{-1}{\bf M}_{4}{\bf A}_{\bm{\theta}^{*}}^{-1}$ (57)
$={\bf A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{1}{\bf A}_{\bm{\theta}^{*}}^{-1}-{\bf A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{4}^{\rm T}({\bf M}_{2}+\rho{\bf M}_{3})^{-1}{\bf M}_{4}{\bf A}_{\bm{\theta}^{*}}^{-1}$ (58)
$={\bf A}_{\bm{\theta}^{*}}^{-1}\left\{{\bf M}_{1}-{\bf M}_{4}^{\rm T}({\bf M}_{2}+\rho{\bf M}_{3})^{-1}{\bf M}_{4}\right\}{\bf A}_{\bm{\theta}^{*}}^{-1}.$ (59)
To connect our protocol with existing methods, we define
$\bm{\Sigma}(\bm{\omega})=\mathbf{A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{1}\mathbf{A}_{\bm{\theta}^{*}}^{-1}+\bm{\omega}^{\rm T}\mathbf{A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{2}\mathbf{A}_{\bm{\theta}^{*}}^{-1}\bm{\omega}^{\rm T}+\bm{\omega}^{\rm T}\rho\mathbf{A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{3}\mathbf{A}_{\bm{\theta}^{*}}\bm{\omega}-2\bm{\omega}^{\rm T}\mathbf{A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{4}\mathbf{A}_{\bm{\theta}^{*}}^{-1}$ (60)
$\bm{\omega}_{tr}^{*}:={\arg\min}_{\bm{\omega}_{tr}}\operatorname{Tr}[\Sigma(\bm{\omega}_{tr})]\text{ where }\bm{\omega}_{tr}=[\omega_{tr},\dots,\omega_{tr}]^{\rm T}\in\mathbb{R}^{K}$ (61)
$\bm{\omega}_{\text{ele}}^{*}:=[\omega_{\text{ele,1}}^{*},\dots,\omega_{\text{ele,K}}^{*}]\in\mathbb{R}^{K}\text{ where }\omega_{\text{ele,k}}^{*}={\arg\min}_{\omega_{\text{ele,k}}}\Sigma(\bm{\omega}_{\text{ele}})_{kk}$ (62)
By the theory of POP-Inf, PPI++, and PPI paper (Miao et al., 2023b, ; Angelopoulos et al., 2023b, ; Angelopoulos et al., 2023a, ), we have
$n^{1/2}(\widehat{\bm{\theta}}_{\texttt{POP-Inf}}-\bm{\theta}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},\bm{\Sigma}({\rm diag}(\bm{\omega}^{*}_{\text{ele}})))$ (63)
$n^{1/2}(\widehat{\bm{\theta}}_{\textnormal{{PPI++}}}-\bm{\theta}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},\bm{\Sigma}({\rm diag}(\bm{\omega}^{*}_{\text{tr}})))$ (64)
$n^{1/2}(\widehat{\bm{\theta}}_{\texttt{PPI}}-\bm{\theta}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}(0,\bm{\Sigma}({\rm diag}({\bf 1})))$ (65)
Based on the proof of theorem 1, we have the
$n^{1/2}(\widehat{\bm{\theta}}(\bm{\omega})-\bm{\theta}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left(0,{\bf A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{1}{\bf A}_{\bm{\theta}^{*}}^{-1}+\bm{\omega}^{\rm T}{\bf A}_{\bm{\eta}}^{-1}{\bf M}_{2}{\bf A}_{\bm{\eta}}^{-1}\bm{\omega}+\rho\bm{\omega}^{\rm T}{\bf A}_{\bm{\eta}}^{-1}{\bf M}_{3}{\bf A}_{\bm{\eta}}\bm{\omega}-2\bm{\omega}^{\rm T}{\bf A}_{\bm{\eta}}^{-1}{\bf M}_{4}{\bf A}_{\bm{\theta}^{*}}^{-1}\right).$ (66)
Plug in $\bm{\omega}$ with ${\rm diag}(\bm{\omega}^{*}_{\text{ele}})A_{\bm{\theta}^{*}}^{-1}{\bf A}_{\bm{\eta}},{\rm diag}(\bm{\omega}^{*}_{\text{tr}})A_{\bm{\theta}^{*}}^{-1}{\bf A}_{\bm{\eta}}$, and ${\rm diag}({\bf 1})A_{\bm{\theta}^{*}}^{-1}{\bf A}_{\bm{\eta}}$, we have
$n^{1/2}(\widehat{\bm{\theta}}({\rm diag}(\bm{\omega}^{*}_{\text{ele}})A_{\bm{\theta}^{*}}^{-1}{\bf A}_{\bm{\eta}})-\bm{\theta}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},\bm{\Sigma}({\rm diag}(\bm{\omega}^{*}_{\text{ele}})))$ (67)
$n^{1/2}(\widehat{\bm{\theta}}({\rm diag}(\bm{\omega}^{*}_{\text{tr}})A_{\bm{\theta}^{*}}^{-1}{\bf A}_{\bm{\eta}})-\bm{\theta}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},\bm{\Sigma}({\rm diag}(\bm{\omega}^{*}_{\text{tr}})))$ (68)
$n^{1/2}(\widehat{\bm{\theta}}({\rm diag}({\bf 1})A_{\bm{\theta}^{*}}^{-1}{\bf A}_{\bm{\eta}})-\bm{\theta}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}(0,\bm{\Sigma}({\rm diag}({\bf 1})))$ (69)
Denote ${\bf C}={\bf A}_{\bm{\theta}^{*}}^{-1}{\bf A}_{\bm{\eta}}$, we have POP-Inf, PPI++, and PPI is asymptotically equivalent to ${\rm diag}(\bm{\omega}^{*}_{\text{ele}}){\bf C},{\rm diag}(\bm{\omega}^{*}_{\text{tr}}){\bf C}$, and ${\rm diag}({\bf 1}){\bf C}$, respectively. This completes the proof. ∎
A.3 Proof of Proposition 2
Proof.
We apply the first-order Taylor expansion to $g(\widehat{\bm{\theta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal U}})$ around $(\bm{\theta}^{*},\bm{\eta},\bm{\eta})$:
$g(\widehat{\bm{\theta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal U}})$ $\approx g(\bm{\theta}^{*},\bm{\eta},\bm{\eta})+\nabla_{\bm{\theta}^{*}}g(\bm{\theta}^{*},\bm{\eta},\bm{\eta})(\widehat{\bm{\theta}}_{{\cal L}}-\bm{\theta}^{*})+\nabla_{\bm{\eta},1}g(\bm{\theta}^{*},\bm{\eta},\bm{\eta})(\widehat{\bm{\eta}}_{{\cal L}}-\bm{\eta})+\nabla_{\bm{\eta},2}g(\bm{\theta}^{*},\bm{\eta},\bm{\eta})(\widehat{\bm{\eta}}_{{\cal U}}-\bm{\eta}),$ (70)
where we used $\nabla_{\bm{\eta},1}$ and $\nabla_{\bm{\eta},2}$ to denote the gradient of $g(\bm{\theta}^{*},\bm{\eta},\bm{\eta})$ w.r.t the first and second $\bm{\eta}$, respectively.
This can be written as a linear function of $(\bm{\theta}^{*},\bm{\eta},\bm{\eta})$:
$g(\widehat{\bm{\theta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal U}})=\bm{\mu}+\bm{\beta}_{1}^{\rm T}\widehat{\bm{\theta}}_{{\cal L}}+\bm{\beta}_{2}^{\rm T}\widehat{\bm{\eta}}_{{\cal L}}+\bm{\beta}_{3}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}},$ (71)
where $\bm{\mu}=g(\bm{\theta}^{*},\bm{\eta},\bm{\eta})-\nabla_{\bm{\theta}^{*}}g(\bm{\theta}^{*},\bm{\eta},\bm{\eta})\bm{\theta}^{*}-2\nabla_{\bm{\eta}}g(\bm{\theta}^{*},\bm{\eta},\bm{\eta})\bm{\eta}$, $\bm{\beta}_{1}=\nabla_{\bm{\theta}^{*}}g(\bm{\theta}^{*},\bm{\eta},\bm{\eta})$, $\bm{\beta}_{2}=\nabla_{\bm{\eta},1}g(\bm{\theta}^{*},\bm{\eta},\bm{\eta})$, $\bm{\beta}_{3}=\nabla_{\bm{\eta},2}g(\bm{\theta}^{*},\bm{\eta},\bm{\eta})$. Since we require $g(\widehat{\bm{\theta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal U}})\stackrel{{\scriptstyle P}}{{\rightarrow}}\bm{\theta}^{*}$, we have $\bm{\mu}={\bf 0}$, $\bm{\beta}_{1}={\bf 1}$, and $\bm{\beta}_{2}+\bm{\beta}_{3}={\bf 0}$. This leads to
$g(\widehat{\bm{\theta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal U}})=\widehat{\bm{\theta}}_{{\cal L}}-\bm{\beta}_{3}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}}),$ (72)
Given
$n^{1/2}\left(\begin{array}[]{c}\widehat{\bm{\theta}}_{{\cal L}}-\bm{\theta}\\ \widehat{\bm{\eta}}_{{\cal L}}-\bm{\eta}\\ \widehat{\bm{\eta}}_{{\cal U}}-\bm{\eta}\\ \end{array}\right)\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left\{\left(\begin{array}[]{c}{\bf 0}_{K}\\ {\bf 0}_{K}\\ {\bf 0}_{K}\end{array}\right),\left(\begin{array}[]{ccc}{\bf V}(\widehat{\bm{\theta}}_{\cal L})&{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf 0}&{\bf 0}&\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U})\end{array}\right)\right\},$ (82)
we have
${\bf V}(g(\widehat{\bm{\theta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal U}}))={\bf V}(\widehat{\bm{\theta}}_{\cal L})+\bm{\beta}_{3}^{\rm T}{\bf V}(\widehat{\bm{\eta}}_{\cal L})\bm{\beta}_{3}+\bm{\beta}_{3}^{\rm T}\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U})\bm{\beta}_{3}-2\bm{\beta}_{3}^{\rm T}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L}),$ (83)
which is a quadratic function of $\bm{\beta}_{3}$ and achieves it minimum when $\bm{\beta}_{3}=({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})=\bm{\omega}_{0}$. This completes the proof. ∎
A.4 Proof of Proposition 3
We state the following definition and lemma that will be used later.
Definition 1 (positive regression dependence on the subset (PRDS)).
For any increasing set $D$, and for each $i\in I_{0},P(\mathbf{X}\in D\mid\left.X_{i}=x\right)$ is non-decreasing in $x$.
Lemma 1.
(Benjamini and Yekutieli,, 2001) If the joint distribution of the test statistics is PRDS on the subset of test statistics corresponding to true null hypotheses, the BH procedure controls the FDR at level less than or equal to $\dfrac{K_{0}}{K}q$.
This lemma is Theorem 1.2 in Benjamini and Yekutieli, (2001), thus the proof is omitted.
We first show that the asymptotic p-value for POP-LM can be derived from the asymptotic distribution of the test statistics. By theorem 1, we have
$n^{1/2}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}-\bm{\theta}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left(0,{\bf V}(\widehat{\bm{\theta}}_{\cal L})-{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})\right),$ (84)
The asymptotic valid p-value can then be calculated by the test statistics
$\dfrac{n^{1/2}\widehat{{\bm{\theta}}}_{\textnormal{{PSPS}}_{k}}}{\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}}\sim\mathcal{N}\left(0,1\right),$ (85)
By the PRDS property on the p-value, applying lemma 1 completes the proof.
A.5 Proof of Proposition 4
This proof is based on (Barber and Candès,, 2019). We state the following definition and lemma that will be used later.
Lemma 2.
(Barber and Candès,, 2019) Suppose that $Z_{1},\ldots,Z_{K}$ are independent with $Z_{k}\sim\operatorname{Bernoulli}\left(\rho_{k}\right)$, and $\min_{k}\rho_{k}\geq\rho>0$. Let $J$ be a stopping time in reverse time with respect to the filtration $\left\{\mathcal{F}_{k}\right\}$, where
$\mathcal{F}_{k}=\sigma\left(\left\{\left(Z_{1}+\cdots+Z_{k}\right),Z_{k+1},\cdots,Z_{d}\right\}\right),$ (86)
with $\sigma(\cdot)$ denoting a $\sigma$-algebra. Then
$\mathbb{E}\left[\frac{1+K}{1+Z_{1}+\cdots+Z_{K}}\right]\leq\rho^{-1}.$ (87)
Define the $S_{k}=\textnormal{sign}(W_{k})\in\{-1,0,1\}$, a random subset of $\mathcal{N}$ as $\mathcal{M}:=\mathcal{N}\backslash\left\{k\in\mathcal{N}:C_{k}=0\right\}=\{k\in\mathcal{N}:S_{k}\neq 0$. Here, we consider the difference statistic
$W_{k}=\left|\hat{\beta}_{k}(\lambda)\right|-\left|\hat{\beta}_{k+K}(\lambda)\right|,k=1,\ldots,K,$ (88)
where $\hat{\beta}_{k}(\lambda)$ and $\hat{\beta}_{k+K}(\lambda)$ is the debiased Lasso coefficient for the true and knockoff variables, respectively.
The FDR is equal to
FDR $=\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:W_{k}\geq\tau^{\textnormal{knockoff}}\right\}\vee 1}\right]$ (89)
$=\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\cdot\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}{\#\left\{k:W_{k}\geq\tau^{\textnormal{knockoff}}\right\}\vee 1}\right]$ (90)
$\leq\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\cdot\dfrac{\#\left\{k:W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}{\#\left\{k:W_{k}\geq\tau^{\textnormal{knockoff}}\right\}\vee 1}\right]$ (91)
$\leq q\cdot\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\right]$ (92)
We reordered the indices of the null $W_{k}^{\prime}s$ so that $\left|W_{(1)}\right|\geq\left|W_{(2)}\right|\geq\ldots$. Let $Z_{k}={\bf 1}_{W_{(k)}}<0$. Let $K$ be the index such that $\left|W_{(1)}\right|\geq\cdots\geq\left|W_{(K)}\right|\geq\tau^{\textnormal{knockoff}}>\left|W_{(K+1)}\right|\geq\cdots$.
For each $k$, we have
$W_{(k)}\geq\tau^{\textnormal{knockoff}}\quad$ $\Leftrightarrow\quad\left|W_{(k)}\right|\geq\tau^{\textnormal{knockoff}}\text{ and }W_{(k)}>0\quad\Leftrightarrow\quad j\leq J\text{ and }Z_{k}=0\text{,}$ (94)
$W_{(k)}\leq-\tau^{\textnormal{knockoff}}\quad$ $\Leftrightarrow\quad\left|W_{(k)}\right|\geq\tau^{\textnormal{knockoff}}\text{and }W_{(k)}<0\quad\Leftrightarrow\quad j\leq J\text{ and }Z_{k}=1$ (95)
Therefore, we have
$\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\right]$ $=\frac{\left(1-Z_{1}\right)+\cdots+\left(1-Z_{k}\right)}{1+Z_{1}+\cdots+Z_{k}}$ (96)
$=\frac{1+K}{1+Z_{1}+\cdots+Z_{k}}-1.$ (97)
By using debiased Lasso coefficient as to construct the $W_{k}$, we have asymptotically $Z_{k}\stackrel{{\scriptstyle\text{ i.i.d. }}}{{\sim}}$ Bernoulli(0.5). With $\rho=0.5$ in lemma 2, we have
$\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\right]\leq 1$ (98)
Therefore,
$\textnormal{FDR}=q\cdot\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\right]\leq q,$ (99)
which completes the proof.
Appendix B Simulation details
Here, we provide the details for our simulation. All our simulation is run in R with version 4.2.1 (2022-06-23) in a macbook air with M1 chip. For all the simulation, the ground truth coefficients are obtained using $5\times 10^{4}$ samples; A pre-trained random forest with 500 trees to grow is obtained from a hold-out data. We bootstrap the labeled data for $200$ times for covariance estimation. All simulations are repeated 1000 times.
  • Mean estimation, Linear regression, and Quantile regression: We simulate the data from $Y_{i}=X_{1i}\beta_{1}+X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}+\epsilon_{i}$, where $X_{i1}$ and $X_{2i}$ are independent simulated from $\mathcal{N}(0,1)$, $\beta_{1}=\sqrt{0.08}$, $\beta_{2}$ is set to be the value such that $X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}$ explains $81\%$ of the outcome variance, and $\epsilon_{i}$ is simulated from a mean zero normal distribution with variance such that $\operatorname{Var}(Y_{i})=1$. We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $1000,2500,5000$, or $10000$.
  • Logistic regression: We simulate the data from $Y_{i}=\mathds{1}(\tilde{Y}_{i}>\text{ median }(\tilde{Y}_{i}))$, where $\tilde{Y}_{i}=X_{1i}\beta_{1}+X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}+\epsilon_{i}$, where $X_{i1}$ and $X_{2i}$ are independent simulated from $\mathcal{N}(0,1)$, $\beta_{1}=\sqrt{0.08}$, $\beta_{2}$ is set to be the value such that $X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}$ explains $81\%$ of the outcome variance, and $\epsilon_{i}$ is simulated from a mean zero normal distribution with variance such that $\operatorname{Var}(\tilde{Y}_{i})=1$. We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $1000,2500,5000$, or $10000$.
  • Instrumental variable (IV) regression: We simulate the data by
    $Z_{i}\sim\mathcal{N}(0,1),$ (100)
    $X_{1i}=0.4Z_{i}+\delta_{i},\delta_{i}\sim\mathcal{N}(0,0.84),$ (101)
    $X_{2i}=0.3Z_{i}+0.8Y_{i}+\gamma_{i},\textnormal{ where }\gamma_{i}\sim\mathcal{N}(0,\tau_{\gamma}),\textnormal{ such that}\operatorname{Var}(X_{2i})=1,$ (102)
    $Y_{i}=\sqrt{0.08}X_{1i}+\epsilon_{i},\textnormal{ where }\epsilon_{i}\sim\mathcal{N}(0,\tau_{\epsilon}),\textnormal{ such that}\operatorname{Var}(Y_{i})=1$ (103)
    We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $1000,2500,5000$, or $10000$. The $Z_{i}$ is used as a instrument for $X_{1i}$.
  • Negative binomial (NB) regression: We simulate the data by
    $X_{1i}\sim\mathcal{N}(0,1),X_{2i}\sim\mathcal{N}(0,1),$ (104)
    $\mu_{i}=\exp(\sqrt{0.3}X_{1i}+0.8X_{2i})$ (105)
    $Y_{i}=\text{NegativeBinomial}\left(s=k,\mu=\mu_{i}\right),\text{where }s\text{ is the dispersion parameter and }\mu\text{ is the rate}.$ (106)
    We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $1000,2500,5000$, or $10000$.
  • Debiased Lasso: We simulate the data by
    $X_{1i},\dots,X_{200i}\sim\mathcal{N}(0,1)$ (107)
    $\theta_{1},\dots,\theta_{15}=\frac{0.9}{\sqrt{15}};\theta_{16},\dots,\theta_{200}=0$ (108)
    $Y_{i}=\sum_{k=1}^{150}X_{ki}\theta_{k}+\epsilon_{i},\epsilon_{i}\sim\mathcal{N}(0,\tau_{\epsilon})\text{ such that }\operatorname{Var}(Y_{i})=1.$ (109)
    We use $X_{1i},\dots,X_{200i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 100 individuals, and unlabeled data with samples size $1500,2000,2500$, or $3000$.
  • Wilcoxon rank-sum test: We simulate the data by
    $X_{1i}\sim\textnormal{Bernoulli}(0.5),X_{2i}\sim\mathcal{N}(0,1)$ (110)
    $Y_{i}=\beta_{1}X_{1i}+\beta_{2}X_{2i}+\beta_{2}X_{2i}^{2}+\epsilon_{i},\epsilon_{i}\sim\mathcal{N}(0,\tau_{\epsilon}),$ (111)
    where $\beta_{1}=\sqrt{0.01}$ to power simulation and $\beta_{1}=0$ for type-I error simulation, $\beta_{2}$ is set to be the value such that $\beta_{2}X_{2i}+\beta_{2}X_{2i}^{2}$ explains $81\%$ of the outcome variance, and $\tau_{\epsilon}$ is set to be the value such that $\operatorname{Var}(Y_{i})=0$.We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $1000,2500,5000$, or $10000$.
  • Benjamini-Hochberg (BH) procedure: We set $K=150$ generate the features independently from $\mathcal{N}(0,\bm{\Sigma})$, where $\bm{\Sigma}$ is a symmetric Toeplitz matrix that has the structure:
    $\bm{\Sigma}=\left[\begin{array}[]{cccc}r^{0}&r^{1}&\ldots&r^{p-1}\\ r^{1}&\ddots&\ldots&r^{p-2}\\ \vdots&\ldots&\ddots&\vdots\\ r^{p-1}&r^{p-2}&\ldots&r^{0}\end{array}\right]$ (116)
    The correlation $r$ is set to be 0.25. We then simulate the outcome $Y_{i}=\sum_{k=1}^{150}X_{ki}\beta_{k}+\epsilon_{i}$, where we randomly sample $15$ $\beta_{k}$ to be 0.15 and let all other remaining $\beta_{k}$ to be 0. $\epsilon_{i}$ is simulated from a mean-zero normal distribution with variance set to the value such that $\operatorname{Var}(Y_{i})=1$. We further generate $Z_{i}=0.9Y_{i}+\sum_{k=1}^{150}X_{ki}\theta_{k}+\gamma_{i}$, where $\theta_{k}=0.15$ for all $k=1,\dots,150$. We use $Z_{i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $5000$.
  • knockoff: We used the same setting as described in the BH procedure above to generate the data. The only difference is that we set $beta_{k}=0.5$ for features associated with the outcome and considered labeled data consisting of 100 individuals, along with unlabeled data comprising a sample size of 1000.